OpenStudy (anonymous):

Holly works in a stable. She's throwing straw into a stall that's 12 feet long by 8 feet wide. If she throws 12 cubic feet of straw each minute, how many minutes will it take until the straw in the stall is 2 feet deep? A. 8 B. 12 C. 16 D. 24

OpenStudy (anonymous):

Volume=Area*Depth=12*8*2 Rate*Time=Volume Time=Volume/Rate=12*8*2/12

OpenStudy (anonymous):

I think it's 8

OpenStudy (anonymous):

Times 2

OpenStudy (anonymous):

it was 16 i got it thank you